A quadrilateral has angles of x degrees, 3x degrees, 90 degrees and (x+20) degrees.
a) Write down an equation in x and solve it.
b) What are the sizes of the angles?

Answer by jim_thompson5910(35256)

A quadrilateral has angles of x degrees, 3x degrees, 90 degrees and (x+20) degrees.
a) Write down an equation in x and solve it.
b) What are the sizes of the angles?
---------------------------------------------------------------------------



a) Write down an equation in x and solve it.


All the angles add to 360 degrees, so,

angle1 + angle2 + angle3 + angle4 = 360

x + 3x + 90 + x+20 = 360

5x + 110 = 360

5x = 360-110

5x = 250

x = 250/5

x = 50


---------------------------------------------------------------------------
b) What are the sizes of the angles?

x = 50
3x = 3*50 = 150
90 = 90
x+20 = 50+20 = 70

The four angles are: 50 degrees, 150 degrees, 90 degrees, 70 degrees